OpenStudy (luigi0210):

Find the derivative:

OpenStudy (luigi0210):

\[\LARGE y=5^{2x}~ln\sqrt{x}\]

OpenStudy (anonymous):

\[ 5^{2x}=25^x \]

OpenStudy (anonymous):

\[ \ln\sqrt x=\frac { \ln x}2 \]

OpenStudy (anonymous):

Use product rule.

OpenStudy (luigi0210):

\[\LARGE (2*5^{2x}~ln\sqrt{x})+(5^{2x}*\frac{1}{\sqrt{x}}*\frac{1}{2}x^{-1/2})\] Uh, I think keeping the 5^2x was a bad idea xD

OpenStudy (luigi0210):

Yea, I probably screwed that up big time.

OpenStudy (luigi0210):

Uhm, \[\LARGE \frac{d}{dx}~5^{2x}=2*5^{2x}*ln5\]?

OpenStudy (anonymous):

Yes

OpenStudy (luigi0210):

\[\LARGE (2*5^{2x}*ln5*ln\sqrt{x})+(5^{2x}*\frac{1}{2x})\] Okay there xD

OpenStudy (anonymous):

You could have done it the easy way...

OpenStudy (luigi0210):

Gotta learn the basic way first I suppose xD